In the reaction below, what volume of PCl5(g) will be produced by the reaction of 18.9 L of chlorine gas with excess phosphorus (P4) if all gas volumes are measured at STP?
P4(g) + Cl2(g) => PCl5(g)(unbalanced)
Enter your answer with two decimal places and no units.

The answer should be 7.56, but I keep getting 6.92.

Balanced P4 + 10Cl2 ----> 4PCl5

1 atm * 18.9L = n (.0821)*(298)

0.7725*(4/10)*22.4

=6.92 , I multiplied it times 22.4 because its asking for volume and everything is in STP.

What am I doing wrong?
Answer #1

Reaction is

P4 + 10 Cl2\rightarrow 4PCl5

Accoding to avogadro's law at STP 1 mole of gas occupy volume 22.414 L then 18.9 L of chlorine = 18.9/22.414 = 0.8432229856 mole

According to reaction 10 mole of clorine produce 4 mole of PCl5 then 0.8432229856 mole of chlorine produce

0.8432229856 \times 4 /10 = 0.3372891942 mole of PCl5

Accoding to avogadro's law at STP 1 mole of gas occupy volume 22.414 L then 0.3372891942 mole of PCl5 occupy volume = 0.3372891942 \times 22.414 = 7.559999 L = 7.56 L